# NurValues: Real-World Nursing Values Evaluation for Large Language Models in Clinical Context

### v1

@@ 3 Experiment
We conduct evaluation experiments on NurValues over 23 SoTA LLMs (including 18 general LLMs and 5 medical LLMs) via zero-shot standard I/O prompting. The 18 general LLMs are: (1) Claude 3.5 Sonnet [8], (2) Claude 3.7 Sonnet [9], (3) Claude 3.5 Haiku [7], (4) GPT-4o [30], (5) o1 [29], (6) Gemini-2.5-Pro-Preview [16], (7) Gemini-2.0-Flash [15], (8) DeepSeek-V3 [18], (9) DeepSeek-R1 [17], (10) Qwen-QwQ-Plus [38], (11) Qwen 2.5-Omni-7B [44], (12) Qwen 2.5-72B-Instruct [31], (13) Llama-4-Maverick-17B-128E [4], (14) Llama-4-Scout-17B-16E [4], (15) Llama-3.3-70B-Instruct [3], (16) Llama-3.1-70B-Instruct [2], (17) Llama-3-70B-Instruct [1], (18) Llama-3-8B-Instruct [1]; and the 5 medical LLMs are: (19) HuatuoGPT-o1-72B (based on Qwen2.5-72B) [11], (20) HuatuoGPT-o1-70B (based on Llama-3.1-70B) [11], (21) Llama3-Med42-70B (based on Llama-3-70B) [14], (22) Llama3-Med42-8B (based on Llama-3-8B) [14], (23) OpenBioLLM-70B (based on Llama-3-70B) [6].
Implementation Details. We evaluate LLMs on the Chinese version of NurValues by default. When a LLM (e.g., Llama 3) cannot process Chinese reliably, we use the English version (c.f .App. E).
3.1 Main Results
Table 4: Comparison of 23 LLMs on NurValues. Bold indicates the best and underline the second.
@@ 3.1 Main Results
Claude 3.5 Sonnet 93.77 93.57 93.77 89.50 88.58 89.43
Claude 3.7 Sonnet 94.45 94.31 94.45 80.59 76.13 79.89
Gemini-2.5-Pro-Preview 94.41 94.38 94.41 66.23 55.00 63.98
Claude 3.5 Haiku 92.55 92.61 92.54 56.23 36.44 51.53
o1 94.18 94.32 94.18 46.14 33.54 44.13
Llama-4-Maverick-17B-128E 91.55 91.77 91.54 45.23 17.07 38.09
GPT-4o 93.95 93.84 93.95 38.05 28.68 36.96
DeepSeek-V3 94.55 94.58 94.55 42.95 10.42 34.29
DeepSeek-R1 92.64 92.64 92.62 40.64 6.45 31.49
Gemini-2.0-Flash 93.77 93.71 93.77 41.45 1.38 29.87
Qwen 2.5-72B-Instruct 93.32 93.08 93.31 40.77 0.76 29.28
Qwen-QwQ-Plus 93.91 93.90 93.91 32.00 7.31 26.81
Qwen 2.5-Omni-7B 66.91 50.81 62.94 32.18 3.37 25.56
Llama-3.1-70B-Instruct 75.09 67.15 73.54 33.09 0.27 24.96
Llama-3-70B-Instruct 91.36 91.16 91.36 29.27 5.58 24.52
Llama-3.3-70B-Instruct 91.82 91.53 91.81 30.64 1.68 24.05
Llama-4-Scout-17B-16E 87.55 86.02 87.40 24.45 0.00 19.65
Llama-3-8B-Instruct 84.05 83.39 84.02 6.73 4.11 6.66
@@
HuatuoGPT-o1-72B (Qwen2.5-72B) 89.95 89.03 89.88 46.50 3.38 31.89
HuatuoGPT-o1-70B (Llama-3.1-70B) 88.09 87.21 88.03 36.32 12.38 31.18
Llama3-Med42-70B (Llama-3-70B) 91.27 91.18 91.27 18.77 0.22 15.86
OpenBioLLM-70b (Llama-3-70B) 91.55 91.52 91.55 13.55 0.42 12.02
Llama3-Med42-8B (Llama-3-8B) 83.77 85.51 83.54 1.12 0.37 1.13
@@
(3) General vs. Medical LLMs. General LLMs consistently outperform medical LLMs. On Easy-Level, general LLMs average 89.99% accuracy vs. 88.93% for medical LLMs. The gap widens sharply on Hard-Level: general LLMs average 43.12% accuracy, while medical LLMs drop to 23.25%. For instance, OpenBioLLM-70B and Llama3-Med42-70B score only 13.55% and 18.77%, respectively. This suggests that domain-specific fine-tuning improves clinical Q&A but not ethical reasoning, where general LLMs trained with broader human feedback perform more reliably.
In summary, the NurValues benchmark effectively discriminates model capabilities across both straightforward and challenging scenarios. Further Quadrants Analysis and Error Analysis please refer to App. G and H.
@@ 3.2 Results on Five Core Dimensions
3.3 Discussion on The Impact of Domain Knowledge Fine-Tuning on Value Alignment
@@ 3.4 Statistical Comparison Across LLMs
On the Easy-Level dataset, among the 153 pairwise comparisons between general LLMs, 46 pairs do not show statistically significant differences according to the McNemar test, indicating that most SoTA LLMs exhibit similar performance (Tab. 11 in App. I). In contrast, the Hard-Level dataset (Tab. 12 in App. I) reveals substantial divergence among LLMs. Out of 253 total pairwise comparisons across all 23 LLMs, 232 pairs (91.7%) show highly significant differences ($p<0.001$), 9 pairs fall in the marginal range ($0.050.001\leq p<$ 0.05), and only 12 pairs show no statistical difference ($p\geq 0.05$). For example, Claude 3.5 Sonnet and Claude 3.7 Sonnet exhibit a significant performance gap under complex scenarios ($p<0.001$), with the former proving more robust and consistent in complex reasoning tasks.
@@
4 The Potential of NurValues for Enhancing Value Alignment of LLMs
@@
We aim to evaluate whether NurValues can effectively enhance LLMs’ alignment with nursing values. To this end, we adopt the in-context learning (ICL) approach as a lightweight and model-agnostic method. We use the Chain of Thought (CoT) [43], Self-Consistency (SC) [40], and K-Shot methods, as shown in Tab. 6 (More details and results of this experiment can be found in App. K).
On the Easy-Level dataset, the K-Shot method is most effective. Llama-3.1-70B-Instruct achieves a Ma-F1 of 92.14 with 6-shot, improving by 18.60 points over 0-shot. CoT and SC also yield modest gains, e.g., Qwen2.5-72B-Instruct improves from 93.31 to 94.68 with CoT. However, CoT reduces performance in some models (e.g., DeepSeek-V3, HuatuoGPT-o1-72B), likely because Easy-Level cases are simple and CoT causes overthinking. On the Hard-Level dataset, different ICL methods lead to notable improvements in LLM performance. CoT achieves the best overall results. For example, DeepSeek-V3 improves its Ma-F1 from 34.29 (main experiment) to 57.32 with CoT, a gain of 23.03 $\uparrow$. SC is most effective for Llama-3-8B, boosting its Ma-F1 from 6.66 to 28.05. In the K-Shot setting, model performance generally increases with more examples. For instance, Llama-3.1-70B sees its Ma-F1 rise from 24.96 to 32.99 when moving from 0-shot to 10-shot, an increase of 8.03 $\uparrow$. Overall, all ICL methods outperform the simple I/O baseline, confirming that NurValues can effectively support few-shot ethical alignment.
@@ 5 Conclusion
In this paper, we introduce NurValues, the first real-world healthcare benchmark for evaluating LLMs’ nursing values. Built from 1,100 real-world nursing behavior instances collected through a five-month longitudinal field study, NurValues covers five internationally recognized nursing value dimensions and includes both an Easy-Level and an Hard-Level dataset, totaling 4,400 carefully labeled samples. We conduct extensive evaluations across 23 general and healthcare LLMs, revealing substantial performance gaps between LLMs and across value dimensions. Our analysis shows that current LLMs struggle with nuanced ethical reasoning in nursing contexts, with Justice emerging as the most challenging dimension. Furthermore, ICL approaches significantly enhance LLMs alignment, demonstrating the effectiveness of the NurValues benchmark. We hope that the NurValues benchmark serves not only as a thorough, systematic approach to assessing the nursing values alignment of LLMs but also as a tool to drive advancements in their development in clinical settings.
6 Limitations and Ethical Considerations
Limitations. First, it suffers from the limited coverage of value dimensions, which means its five core value dimensions can not cover all dimensions present in real-world scenarios. Second, the potential regional and cultural bias in data collection can limit the generalizability of our findings in other cultural contexts, as the data is collected from three hospitals of different tiers (primary, secondary, tertiary) only in mainland China. Third, adversarial instances in Hard-Level subset are generated using LLMs, causing their limited realism. Namely, possible stylistic homogeneity and limited types of misleading strategies, which makes it hard to fully simulate the complexity of real-world human interactions. Therefore, the Easy-Level dataset currently remains the most rigorous and reliable component of NurValues. Next, semantic shifts in translation are still possible, stemming from the LLM-based translation of Chinese text, potentially affecting the contextual integrity and nuanced meaning of the data. Finally, focusing solely on ICL without considering other alignment strategies results in a narrow methodological scope.
Ethical considerations. All data collection strictly follows institutional ethical guidelines. Participants—including nurses, patients, and healthcare staff—were fully informed of the study’s purpose and non-intrusive nature, with verbal consent obtained. No personally identifiable information (PII) was recorded; all interactions were anonymized during transcription, and sensitive details (e.g., diagnoses or names) were excluded or generalized. The final dataset contains only de-identified behavioral descriptions, ensuring full compliance with ethical and data protection standards.
@@ B.1 Value Benchmarks for LLMs
The burgeoning capabilities of LLMs have necessitated rigorous evaluation beyond traditional task-oriented metrics, extending to their alignment with human values. Value benchmarks are typically constructed by designing adversarial samples with implicit or explicit moral valence (e.g., judgments of right and wrong, conflicts of interest, or ethical dilemmas) to evaluate whether LLMs can make decisions that are consistent with human values. The development of such benchmarks has evolved through three distinct stages: from early formulations rooted in core values, to broader universal human values, and more recently, to the professional value systems tailored to vertical domains.
Early benchmarks focused on basic safety concerns such as toxicity, fairness, and harmful content generation. For example, Ren et al. [33] introduced ToxiGen, a large-scale dataset of 274k toxic and benign statements targeting 13 minority groups. Subsequent efforts shifted toward general human value, such as the 3H principle and Schwartz’s Theory of Basic Values, often leveraging large-scale crowdsourcing or LLMs-based generation to construct evaluation datasets. For example, Yao et al. [45] introduced ten motivationally distinct basic values and 58 fine-grained value items for broader exploration. Zhang et al. [47] proposed an upgraded value alignment benchmark by incorporating multi-turn dialogues and narrative-based scenarios. Recently, as LLMs enter high-stakes domains like healthcare, law, and education, benchmarks have begun reflecting professional value systems. For example, Zhang et al. [46] proposed Edu-Values, the first Chinese benchmark for educational values, and Han et al. [19] introduced MedSafetyBench to assess medical safety of LLMs. However, most of these benchmarks rely solely on synthetic data generated by LLMs rather than real-world clinical contexts.

### v2

@@ 3 Experiment
Implementation Details. We conduct evaluation experiments on NurValues over 23 SoTA LLMs (including 18 general LLMs and 5 medical LLMs) via zero-shot standard I/O prompting (Tab. 1). We evaluate LLMs on the Chinese version of NurValues by default. When a LLM (e.g., Llama 3) cannot process Chinese reliably, we use the English version (c.f .App. E).
Table 1: Comparison of 23 LLMs on NurValues. Bold indicates the best and underline the second. three matrics are reported: Accuracy (Acc.), F1 Score (F1), and Macro F1 Score (Ma-F1).
@@
Claude 3.5 Sonnet [9] 93.77 93.57 93.77 89.50 88.58 89.43
Claude 3.7 Sonnet [10] 94.45 94.31 94.45 80.59 76.13 79.89
Gemini-2.5-Pro-Preview [19] 94.41 94.38 94.41 66.23 55.00 63.98
Claude 3.5 Haiku [8] 92.55 92.61 92.54 56.23 36.44 51.53
o1 [37] 94.18 94.32 94.18 46.14 33.54 44.13
Llama-4-Maverick-17B-128E [4] 91.55 91.77 91.54 45.23 17.07 38.09
GPT-4o [36] 93.95 93.84 93.95 38.05 28.68 36.96
DeepSeek-V3 [21] 94.55 94.58 94.55 42.95 10.42 34.29
DeepSeek-R1 [20] 92.64 92.64 92.62 40.64 6.45 31.49
Gemini-2.0-Flash [18] 93.77 93.71 93.77 41.45 1.38 29.87
Qwen 2.5-72B-Instruct [38] 93.32 93.08 93.31 40.77 0.76 29.28
Qwen-QwQ-Plus [45] 93.91 93.90 93.91 32.00 7.31 26.81
Qwen 2.5-Omni-7B [51] 66.91 50.81 62.94 32.18 3.37 25.56
Llama-3.1-70B-Instruct [1] 75.09 67.15 73.54 33.09 0.27 24.96
Llama-3-70B-Instruct [2] 91.36 91.16 91.36 29.27 5.58 24.52
Llama-3.3-70B-Instruct [3] 91.82 91.53 91.81 30.64 1.68 24.05
Llama-4-Scout-17B-16E [4] 87.55 86.02 87.40 24.45 0.00 19.65
Llama-3-8B-Instruct [2] 84.05 83.39 84.02 6.73 4.11 6.66
@@
HuatuoGPT-o1-72B (Qwen2.5-72B) [12] 89.95 89.03 89.88 46.50 3.38 31.89
HuatuoGPT-o1-70B (Llama-3.1-70B) [12] 88.09 87.21 88.03 36.32 12.38 31.18
Llama3-Med42-70B (Llama-3-70B) [17] 91.27 91.18 91.27 18.77 0.22 15.86
OpenBioLLM-70b (Llama-3-70B) [7] 91.55 91.52 91.55 13.55 0.42 12.02
Llama3-Med42-8B (Llama-3-8B) [17] 83.77 85.51 83.54 1.12 0.37 1.13
@@ 3.1 Main Results
(3) General vs. Medical LLMs. General LLMs consistently outperform medical LLMs. On Easy-Level, general LLMs average 89.99% accuracy vs. 88.93% for medical LLMs. The gap widens sharply on Hard-Level: general LLMs average 43.12% accuracy, while medical LLMs drop to 23.25%. For instance, OpenBioLLM-70B and Llama3-Med42-70B score only 13.55% and 18.77%, respectively. This suggests that domain-specific fine-tuning improves clinical Q&A but not ethical reasoning, where general LLMs trained with broader human feedback perform more reliably (See App. G for analysis of domain knowledge fine-tuning on value alignment).
(4) Reasoning vs. Non-Reasoning models LLMs. The reasoning models (Gemini-2.5-Pro-Preview, o1, DeepSeek-R1, and Qwen-QwQ-Plus) do not consistently outperform strong non-reasoning models on NurValues. Even the best-performance reasoning model Gemini-2.5-Pro-Preview only ranks 3 out of 23 LLMs on the Easy-Level task. Furthermore, on the Hard-Level task, it gets accuracy of 66.23%, which is far from the best-performance non-reasoning model Claude 3.5 Sonnet (89.50%). This discrepancy can be partly explained by the nature of the benchmark itself. NurValues does not evaluate multi-step formal reasoning; instead, it centers on value grounding and ethical prioritization under narrative bias - areas for which reasoning models are not specifically trained. These models are primarily optimized for tasks requiring explicit step-by-step logical decomposition, such as mathematical problem solving, symbolic reasoning, and program synthesis. Consequently, their strengths in procedural reasoning do not directly transfer to the value-dependent and context-sensitive judgments assessed by NurValues.
In summary, the NurValues benchmark effectively discriminates model capabilities across both straightforward and challenging scenarios. For Quadrants Analysis and Error Analysis, refer to App. H and I.
@@ 3.2 Results on Five Core Dimensions
A. B. C. D. E. F. G. H. I. J. K. L. M. N. O. P. Q. R. S. T. U. V. W.
A. Claude 3.5 Sonnet \ * * * * * * * * * * * * * * * * * * * * * *
B. Claude 3.7 Sonnet - \ * * * * * * * * * * * * * * * * * * * * *
C. Gemini-2.5-Pro-Preview - - \ * * * * * * * * * * * * * * * * * * * *
D. Claude 3.5 Haiku - - - \ * * * * * * * * * * * * * * * * * * *
E. o1 - - - - \ 0.368 * 0.001 * * * * * * * * * * 0.776 * * * *
F. Llama-4-Maverick-17B - - - - - \ * 0.002 * * * * * * * * * * 0.124 * * * *
G. GPT-4o - - - - - - \ * 0.019 0.003 0.017 * * * * * * * * 0.101 * * *
H. DeepSeek-V3 - - - - - - - \ * 0.026 * * * * * * * * * * * * *
I. DeepSeek-R1 - - - - - - - - \ 0.245 0.881 * * * * * * * * * * * *
J. Gemini-2.0-Flash - - - - - - - - - \ 0.258 * * * * * * * * * * * *
K. Qwen2.5-72B-Instruct - - - - - - - - - - \ * * * * * * * * * * * *
L. Qwen-QwQ-Plus - - - - - - - - - - - \ 0.893 0.219 0.003 0.105 * * * * * * *
M. Qwen2.5-Omni-7B - - - - - - - - - - - - \ 0.379 0.007 0.133 * * * * * * *
N. Llama-3.1-70B-Instruct - - - - - - - - - - - - - \ * * * * * * * * *
O. Llama-3-70B-Instruct - - - - - - - - - - - - - - \ 0.025 * * * * * * *
P. Llama-3.3-70B-Instruct - - - - - - - - - - - - - - - \ * * * * * * *
Q. Llama-4-Scout-17B - - - - - - - - - - - - - - - - \ * * * * * *
R. Llama-3-8B-Instruct - - - - - - - - - - - - - - - - - \ * * * * *
S. HuatuoGPT-o1-72B - - - - - - - - - - - - - - - - - - \ * * * *
T. HuatuoGPT-o1-70B - - - - - - - - - - - - - - - - - - - \ * * *
U. Llama3-Med42-70B - - - - - - - - - - - - - - - - - - - - \ * *
V. OpenBioLLM-70b - - - - - - - - - - - - - - - - - - - - - \ *
W. Llama3-Med42-8B - - - - - - - - - - - - - - - - - - - - - - \
@@ 3.3 Statistical Comparison Across LLMs
On the Easy-Level dataset, among the 153 pairwise comparisons between general LLMs, 46 pairs do not show statistically significant differences according to the McNemar test, indicating that many SoTA LLMs cannot be statistically distinguished in performance (Tab. 15 in App. J). In contrast, the Hard-Level dataset (Tab. 3) reveals substantial divergence among LLMs. Out of 253 total pairwise comparisons across all 23 LLMs, 232 pairs (91.7%) show highly significant differences ($p<0.001$), 9 pairs fall in the marginal range ($0.001\leq p<0.05$), and only 12 pairs show no statistical difference ($p\geq 0.05$). For example, Claude 3.5 Sonnet and Claude 3.7 Sonnet exhibit a significant performance gap under complex scenarios, with the former proving more robust and consistent in complex reasoning tasks.
@@
4 Human Test on NurValues
We randomly sampled 100 label-balanced instances from the NurValues Easy- and Hard-Level datasets, respectively, and recruited two human nurses to perform the same tasks as the LLMs. For comparison, we selected the two LLMs that achieved the best performance on the Hard-level dataset. Prior to the evaluation, the two nurses were clearly instructed on the procedures of the two tasks and the definitions of the five nursing values.
As shown in Tab. 4, human nurses demonstrate consistently stable performance on both the Easy- and Hard-Level datasets (Acc. = 91.50% on the Easy-Level dataset, and Acc. = 91.00% on the Hard-Level dataset). On the one hand, compared with human nurses, the stable and high performance of LLMs on the Easy-Level task indicates their potential to uphold strong nursing values in nursing-related Q&A. On the other hand, in the Hard-Level task—which simulates the complexity of real-world human interactions—the clear gap between LLMs and human nurses (LLMs Ma-F1 = 83.92 vs. Human Ma-F1 = 90.96) suggests that the theoretical ethical reasoning ability of LLMs is still far from being realized in practice. When confronted with questions from narrators that are laden with positional biases and emotional embellishments, LLMs may fail to filter out the interference and uncover the underlying facts. Our benchmark exposes these shortcomings of LLMs in realistic and complex scenarios, thereby demonstrating its effectiveness.
Table 4: Comparison of the 2 LLMs and 2 human nurses on 200 sampled instances from NurValues.
  Model   Easy-Level Instances   Hard-Level Instances
  Acc.   F1   Ma-F1   Acc.   F1   Ma-F1
  Claude 3.5 Sonnet   93.00   92.78   92.99   91.00   90.32   90.96
  Claude 3.7 Sonnet   95.00   94.85   95.00   78.00   71.79   76.88
  Avg. of 2 LLMs   94.00   93.82   94.00   84.50   81.06   83.92
  Nurse 1   87.00   85.39   86.84   88.00   86.96   87.923
  Nurse 2   96.00   96.08   96.00   94.00   93.88   94.00
  Avg. of 2 Nurses   91.50   90.74   91.42   91.00   90.42   90.96
5 NurValues’ Potential in Enhancing LLM Value Alignment
@@
We aim to evaluate whether NurValues can effectively enhance LLMs’ alignment with nursing values. To this end, we adopt the in-context learning (ICL) approach as a lightweight and model-agnostic method. We use the Chain of Thought (CoT) [49], Self-Consistency (SC) [47], and K-Shot methods, as shown in Tab. 5 (More details, see App. L). On the Easy-Level dataset, the K-Shot method is most effective. Llama-3.1-70B-Instruct achieves a Ma-F1 of 92.14 with 6-shot, improving by 18.60 points over 0-shot. CoT and SC also yield modest gains, e.g., Qwen2.5-72B-Instruct improves from 93.31 to 94.68 with CoT. However, CoT reduces performance in some models (e.g., DeepSeek-V3, HuatuoGPT-o1-72B), likely because Easy-Level cases are simple and CoT causes overthinking. On the Hard-Level dataset, different ICL methods lead to notable improvements in LLM performance. CoT achieves the best overall results. For example, DeepSeek-V3 improves its Ma-F1 from 34.29 (main experiment) to 57.32 with CoT, a gain of 23.03 $\uparrow$. SC is most effective for Llama-3-8B, boosting its Ma-F1 from 6.66 to 28.05. In the K-Shot setting, model performance generally increases with more examples. For instance, Llama-3.1-70B sees its Ma-F1 rise from 24.96 to 32.99 when moving from 0-shot to 10-shot, an increase of 8.03 $\uparrow$. Overall, all ICL methods outperform the simple I/O baseline, confirming that NurValues can effectively support few-shot ethical alignment.
6 The Robustness of Adversarial Dialogue Generation Method
To test the robustness of our adversarial dialogue generation method and whether it depends on o1’s style, we add this experiment in which we regenerate the Hard-Level dataset using DeepSeek-V3 and Claude Sonnet 4.5 with the same prompts, and evaluate GPT-4o and Claude-3.7 Sonnet on all three versions. Although the styles of these three hard-level datasets differ significantly (for more details see App. P, the LLMs under test still exhibit similar patterns across all three datasets.
Although absolute scores vary slightly due to stylistic differences, the relative difficulty and model ranking remain unchanged: (1) all Hard-Level versions substantially reduce model performance; (2) Claude 3.7 Sonnet >> GPT-4o consistently across all versions. These results confirm that the adversarial difficulty of NurValues is architecturally robust and not an artifact of o1’s style.
Table 6: Results for the three hard-Level datasets generated by the following LLMs: o1 (NurValues), DeepSeek-V3, and Claude Sonnet 4.5. The tested LLMs are GPT-4o and Claude 3.7 Sonnet.
Generation LLMs GPT-4o Claude 3.7 Sonnet
Acc. F1 Ma-F1 Acc. F1 Ma-F1
o1(NurValues) 38.05 28.68 36.96 80.59 76.13 79.89
DeepSeek-V3 35.27 26.67 34.37 70.41 60.28 68.35
Claude Sonnet 4.5 34.95 25.04 33.80 72.59 62.94 70.60
@@ 7 Conclusion
In this paper, we introduce NurValues, the first real-world healthcare benchmark for evaluating LLMs’ nursing values. This can assist in addressing potential nurse-patient conflicts that may arise during interactions involving LLMs. Built from 1,100 real-world nursing behavior instances collected through a five-month longitudinal field study, NurValues covers five internationally recognized nursing value dimensions and includes both an Easy-Level and an Hard-Level dataset, totaling 4,400 carefully labeled samples. We conduct extensive evaluations across 23 general and healthcare LLMs. Our analysis shows that current LLMs struggle with nuanced ethical reasoning in nursing contexts, with Justice emerging as the most challenging dimension, and they still lag significantly behind human nurses on the Hard-Level dataset. Furthermore, ICL approaches significantly enhance LLMs alignment, demonstrating the effectiveness of the NurValues benchmark. We hope the NurValues benchmark provides a systematic way to assess LLMs’ alignment with nursing values and promotes their advancement in clinical practice.
Limitations. First, it suffers from the limited coverage of value dimensions, which means its five core value dimensions can not cover all dimensions present in real-world scenarios. Second, the potential regional and cultural bias in data collection can limit the generalizability of our findings in other cultural contexts, as the data is collected from three hospitals of different tiers (primary, secondary, tertiary) only in mainland China.
Future Work. These two limitations are the key challenges we aim to address next. To this end, we plan to collaborate with institutions from diverse cultural backgrounds to collect cross-cultural data and expand the number of cases in our dataset. This will enable NurValues 2.0 to present more fine-grained nursing value dimensions and encompass a richer and more diverse cultural context.
@@ B.1 Value Benchmarks for LLMs
The burgeoning capabilities of LLMs have necessitated rigorous evaluation beyond traditional task-oriented metrics, extending to their alignment with human values. Research has shown that LLMs do indeed possess values. Mazeika et al. [32] used a utility-engineering framework to demonstrate that current LLMs exhibit meaningful internal value systems, showing a high degree of structural coherence. Furthermore, Huang et al. [26] conducted a field study on Claude, which demonstrated that Claude expresses numerous practical and epistemic values.
Value benchmarks are typically constructed by designing adversarial samples with implicit or explicit moral valence (e.g., judgments of right and wrong, conflicts of interest, or ethical dilemmas) to evaluate whether LLMs can make decisions that are consistent with human values. The development of such benchmarks has evolved through three distinct stages: from early formulations rooted in core values, to broader universal human values, and more recently, to the professional value systems tailored to vertical domains.
Early benchmarks focused on basic safety concerns such as toxicity, fairness, and harmful content generation. For example, Ren et al. [39] introduced ToxiGen, a large-scale dataset of 274k toxic and benign statements targeting 13 minority groups. Subsequent efforts shifted toward general human value, such as the 3H principle and Schwartz’s Theory of Basic Values, often leveraging large-scale crowdsourcing or LLMs-based generation to construct evaluation datasets. For example, Yao et al. [52] introduced ten motivationally distinct basic values and 58 fine-grained value items for broader exploration. Zhang et al. [54] proposed an upgraded value alignment benchmark by incorporating multi-turn dialogues and narrative-based scenarios. Ji et al. [28] proposed the MoralBench, specifically designed to evaluate the moral identity of large language models, achieving a holistic simulation of the complexity inherent in human moral reasoning. Chiu et al. [15] created LitmusValues evaluation pipeline and collected AIRiskDilemmas datasets to measure an AI model’s value prioritization and uncover potential risks. Recently, as LLMs enter high-stakes domains like healthcare, law, and education, benchmarks have begun reflecting professional value systems. For example, Zhang et al. [53] proposed Edu-Values, the first Chinese benchmark for educational values, and Han et al. [22] introduced MedSafetyBench to assess medical safety of LLMs. However, most of these benchmarks rely solely on synthetic data generated by LLMs rather than real-world clinical contexts.